OpenStudy (anonymous):

verify the following using trigonometric identities 1-sin^2/cos = cos

OpenStudy (anonymous):

start with \[\cos^2(x)+\sin^2(x)=1\] then \[\cos^2(x)=1-\sin^2(x)\] replace in the numerator, and you are done in one step